OpenStudy (anonymous):

Electrolysis problem anyone...? Given: What current flow is required for this electroplating if the bumper is to be plated in 15.0s ? (A: (5.5*10^6 Coulombs/15s = 3.6*10^5 Amperes) Can't figure out the next problem... If the external source has emf + 5.0V & the electrolytic cell is 66% efficient, how much electrical energy is expended to electroplate the bumper? (A: 12 kWh. Help appreciated with this one, thanks..)

OpenStudy (anonymous):

This is what I did, but I didn't get 12 kWh... Power: Volts X amps => (5.00V*3.6*10^5 A = 1800000 watts) (1800000 watts*10^-3) = 1800 kilowatts Energy (kWh): power X time (in hours) => ( 1800 kW) * (4.167*10^-3 hours) = 7.5 kWh Then since 66% of the cell is efficient.... => .66 * 7.5kWh = 4.95 kWh. But this isn't correct... Could anyone please help me figure this problem out?

OpenStudy (zpupster):

(5.5*10^6)(100/65) = 8.333E6 (8.33E6)(5)(1/3.6E6)= 11.56 kwh

OpenStudy (anonymous):

THANK YOU, I was lost on this problem for a while now. You're the best =]